# ToddlerBot: Open-Source ML-Compatible Humanoid Platform for Loco-Manipulation

### v2

@@ VII Conclusions
In conclusion, we demonstrate that ToddlerBot is ML-compatible, capable, and reproducible through a series of tests and loco-manipulation tasks. While humanoid research is often associated with locomotion, ToddlerBot extends beyond this to support full-body manipulation, character animation, human-robot interaction, and various ML applications, making it a versatile research platform. With ToddlerBot fully open-source, we hope to empower researchers to explore new directions in humanoid research and encourage open collaboration in the community.
@@ Acknowledgments
The authors would like to express their gratitude to Kaizhe Hu for assembling the second instance of ToddlerBot and assisting with keyframe animation and demo recording. We also extend our thanks to Huy Ha, Yen-Jen Wang, Pei Xu, and Yifan Hou for their insightful discussions on locomotion, and to Sirui Chen, Chen Wang, and Yunfan Jiang for valuable input on manipulation policy deployment. We are grateful to Albert Wu for his guidance on mathematical formulation and notation. Additionally, we thank João Pedro Araújo for his assistance with the motion capture system. Finally, we appreciate the helpful discussions from all members of TML and REALab. This work was supported by National Science Foundation NSF-FRR-2153854, NSF-2143601, NSF-2037101, Sloan Fellowship, Stanford Institute for Human-Centered Artificial Intelligence, and Stanford Wu Tsai Human Performance Alliance.

### v3

@@ 6 Conclusions
In conclusion, we demonstrate that ToddlerBot is ML-compatible, capable, and reproducible through a series of tests and loco-manipulation tasks. ToddlerBot extends beyond locomotion and supports full-body manipulation, character animation, human-robot interaction, and a range of ML applications, making it a versatile research platform. To bridge the sim-to-real gap, we develop a comprehensive sysID pipeline to create a high-fidelity digital twin, alongside a whole-body teleoperation system to collect loco-manipulation data. Being fully open-source, ToddlerBot empowers researchers to explore new research directions and fosters open collaboration across the community.
7 Limitations and Future Work
ToddlerBot’s performance in more agile tasks is constrained by the off-the-shelf motors’ max speed, max torque, and communication speed. Should there be stronger motor choices, the performance can still improve. Rather than achieving superhuman capabilities, ToddlerBot aligns more closely with average human performance in loco-manipulation tasks. In addition, the current actuation model does not consider motor temperature and is less accurate when close to the performance limit. The miniature scale limits its interaction with human-sized objects, though this does not hinder research if appropriately sized objects are used. Compared to metal shells, 3D-printed parts are more likely to break after the event of impact. Despite 3D printing being fast, repairs still require time.
To overcome these limitations, we plan to develop customized communication boards to increase control frequency and improve our actuation model to maximize usable envelope, especially when the motor heats up. We will refine ToddlerBot’s design to improve structural strength, as we believe that design has a greater impact than material choice. We also aim to improve sensing capabilities, including stereo vision for depth perception, additional IMUs for improved state estimation, and tactile sensors for richer manipulation feedback.
@@ Acknowledgments
The authors would like to express their gratitude to Kaizhe Hu for assembling the second instance of ToddlerBot and assisting with keyframe animation and demo recording. We also extend our thanks to Huy Ha, Yen-Jen Wang, Pei Xu, and Yifan Hou for their insightful discussions on locomotion, and to Sirui Chen, Chen Wang, and Yunfan Jiang for valuable input on manipulation policy deployment. We are grateful to Albert Wu for his guidance on mathematical formulation and notation. Additionally, we thank João Pedro Araújo for his assistance with the motion capture system. Finally, we appreciate the helpful discussions from all members of TML and REALab. This work was supported by National Science Foundation NSF-FRR-2153854, NSF-2143601, NSF-2037101, NSF-2132519, Sloan Fellowship, Stanford Institute for Human-Centered Artificial Intelligence, and Stanford Wu Tsai Human Performance Alliance.